# A voice for minorities: diversity in approval-based committee elections under incomplete or inaccurate information

### v1

@@ 4.2 Upper bound for a setting with matroid constraints
For any $\delta>0$, $\gamma\in(0,1)$, $m\geq t>k$, and some universal constant $C_{2}$, with probability at least $1-\delta$, Algorithm 4, with $\beta=C_{2}\frac{1-\gamma}{\gamma k\log k}$, is $(1-1/\operatorname{\text{e}})\gamma$-approximate for Max Cover when using
$\xi\alpha_{k}\cdot\left(\frac{\gamma k\log k}{C_{2}(1-\gamma)}\right)^{3}\cdot t\left\lceil\frac{m-k}{t-k}\right\rceil\cdot\left\lceil\frac{(2-2/\operatorname{\text{e}})^{2}}{2}\left(\frac{2\xi}{\xi-1}\right)^{2}\log\left(\frac{2\cdot(m-k)k\cdot\xi\alpha_{k}\cdot\gamma k\log k}{C_{2}(1-\gamma)\delta}\right)\right\rceil$
queries. Asymptotically in $k,m,\gamma$ and $\delta$, this is in
$\mathcal{O}\left(\left(\frac{\gamma k\log k}{1-\gamma}\right)^{3}m\log\left(\frac{m\gamma k^{2}\log k}{\delta(1-\gamma)}\right)\right).$
Example 4.11.
The number of queries increases in $\xi$ when $\xi\geq 3$, and has a negative asymptote at $\xi=1$, so regardless of the values of the other parameters, we attain a minimum for some $\xi\in[1,3]$. Fixing $\xi=3$, for $m=1000,\,\delta=0.05,\,k=8$ and $t=20$ the upper bound of Theorem 4.10 gives $9.52180\cdot 10^{11}$ queries; quite a lot.
5 Inaccurate information
In the inaccurate information setting, voters do hand in complete ballots, but the voters are inaccurate in reporting their true approvals: each time we query a voter $i$ about a candidate $j$ (so, posing query $(i,j)$), the response is corrupted with probability $p$. We model this by adding, to the original complete information approval profile $A$, a random variable $X\sim\text{Bernoulli}(p)$, with $p\in(0,\frac{1}{2})$. Posing query $(i,j)$, we also draw from $\text{Bernoulli}(p)$, and the result determines whether we see a truthful response (draw 0) or not (draw 1). Draws from $X$ are independent between voters, between candidates, and between consecutive draws for the same query. If $p=0$, we would have accurate information, and with $nm$ queries we would have all, accurate, information. We would then be sure to find a $(1-1/\operatorname{\text{e}})$-approximate solution (e.g. using the greedy algorithm). Now $p>0$, so without adapting the algorithms, their performance will drop. We here present theoretical results on the amount of queries required to still obtain a $(1-1/\operatorname{\text{e}})$-approximate solution with high probability.
When answers may be corrupted, we can simply pose every question multiple times to compensate for uncertainty. That is the strategy we use to obtain the upper bound. We can adapt Algorithm 1 (attaining Algorithm 5 , see Appendix) or Algorithm 3; it doesn’t matter, since the new total amount of queries only depends, besides on $p$ and $\delta$, on the original amount of queries, which would be $nm$ for both algorithms. This also means we immediately have an upper bound for the problem over a matroid constraint.
Theorem 5.1.
In the query model with inaccurate query responses and parameter $p\in(0,\frac{1}{2})$, for any $\delta>0$, $n,m\in\mathbb{N}$, we attain a $(1-1/\operatorname{\text{e}})$-approximate solution to Max Cover with probability at least $1-\delta$ using $nm\cdot U$ queries, where $U=\lceil-2\frac{\log((nm)/\delta)}{\log(4p(1-p))}\rceil$. This gives a query complexity in $\mathcal{O}(nm\log(nm/\delta))$.
We prove the lower bound using an information theoretic lower bound from multi-armed bandit theory. The proof disregards the possibility of matroid constraints, but since the original problem is a special case of the problem over a matroid constraint, the lower bound holds for both settings.
Theorem 5.2.
@@ 5 Inaccurate information
In the query model with inaccurate query responses, for any $\delta\in(0,1),\,n,m\in\mathbb{N}$, any algorithm that outputs an ($1-1/\operatorname{\text{e}}$)-approximate solution to Max Cover with probability at least $1-\delta$, makes $\Omega(nm\log(1/\delta))$ queries in expectation.
6 Experiments
Our experimental section is motivated by three questions. First, although our theoretical analysis (Theorems 4.4 and 4.10) provides good asymptotic bounds on the query complexity of both Greedy and Local Search-$\beta$ (Algorithms 2 and 4), large constant factors are a big influence for relatively small instances. For both theorems, the number of voters $\ell$ that needs to be sampled for each query set in order to arrive at the stated approximation ratio with high probability, is very large. This is not only impractical (each voter would need to reply more than $n$ times) but also makes the computations extremely slow. However, as the theoretical bounds are worst-case, we might be able to attain good results even on smaller instances. Second, Theorem 5.1 provides a good asymptotic bound on the number of queries required to compensate inaccurate voting behavior, requiring a number $U\in\mathcal{O}(nm/\delta)$ of repetitions for each query. We are interested in the experimental performance of our algorithms when we ignore this advice (that is, take $U=1$). Third and finally, we are interested in the performance of our algorithms when we combine incomplete and inaccurate information. We did not study theoretically in this paper, but a naive approach for an upper bound would be to combine both separate bounds (by multiplication). What can we say about this experimentally?
We investigate these questions by running both our querying algorithms (Algorithms 2 and 4) and their exact versions (Algorithms 1 and 3) on 18 open-use datasets provided by Polis  [^5] https://github.com/compdemocracy/openData, using $p=0$ and $p=\frac{1}{10}$. We used a committee size $k=8$ in all experiments (changing $k$ resulted in very similar results).
@@ 6 Experiments
(a) Mean and standard deviation of the Max Cover score of Greedy (Algorithm 1), Local Search-$\beta$ (Algorithm 3) and Approval Voting on the 18 Polis datasets at each iteration of the algorithms for $k=8$. Since the score of Local Search-$\beta$ depends on the randomly sampled initial committee, we ran 50 trials of that algorithm.
(b) Mean and standard deviation of the Max Cover score of Greedy-queries (Algorithm 2) and Local Search-queries (Algorithm 4) on the 18 Polis datasets at each iteration of the algorithms for $k=8$. We ran 50 trials of both algorithms, as the performance depends on which voters are randomly sampled, as well as on the initial committee, for Local Search-queries. The score is expressed as a fraction of the score of the exact version of the algorithms on the given data set.
Figure 2: Performance of our algorithms, mean and standard deviation taken over 18 Polis datasets
We start with the full information case. In Figure 2(a) we show the Max Cover score, obtained by Greedy (Algorithm 1) and Local Search-$\beta$ (Algorithm 3), per iteration. We took $\beta$ as in Theorem 4.10 (and will write Local Search going forward). We also show the Max Cover score that Approval Voting (selecting the $k$ most popular statements) would reach. Note that Greedy and Approval Voting each have a committee of size $i$ in iteration $i$, whereas Local Search starts with a random committee of size $k$, swapping one committee member each iteration. This explains the differences in scores for low iteration numbers. After $k=8$ iterations, we see that the score is confidently above $1-1/\operatorname{\text{e}}\approx 0.63$, indicating that the approximation ratio is achieved, even when the optimal score is 1. We also see that Greedy and Local Search perform better than Approval Voting, as expected.
@@
Taking the scores of the Greedy and Local Search from Figure 2(a) as our baseline, we want to understand how close we can get to these scores when using only few queries, and/or when information is inaccurate. Recall that $t$ denotes the number of statements contained in a single query set. Taking $t=20$, we write $M$ for the expected number of query sets of size $t$ that each voter is given. We then configure Local Search-queries (Algorithm 4), and Greedy-queries (Algorithm 2) so that $M=1,2,3,4,5$. This is far lower than what would be theoretically required. We do this for both $p=0$ and $p=\frac{1}{10}$. Figure 2(b) shows the performance of both algorithms relative to their exact versions.
@@
We see that both algorithms with accurate responses ($p=0$) obtain a score very close ($0.9-0.95$) to the exact versions of the algorithms, even when $M=1$. For the same number of queries, the local search algorithm seems to perform slightly closer to its exact version than the greedy algorithm does. Note that the exact versions of both algorithms obtained an average Max Cover score of around $0.73$. Multiplying by $0.9$ still yields a score above the worst-case approximation ratio of $1-1/\operatorname{\text{e}}\approx 0.63$, even when the optimal solution would have a Max Cover score of 1. Example 6.1 highlights the exceptional performance witnessed in the experiments for $p=0$.
Example 6.1.
@@
The dataset vtaiwan.uberx has $m=197$ and $n=1921$. In Figure 2(b), we see that, for $M=5$, the Max Cover score attained by Local Search-queries, on average across 18 data sets (and 20 runs per set) is approximately 0.95 of the score attained by Local Search. So, for our example, let’s take $\gamma=0.95$. Then, for $k=8$ and $t=20$, taking $\delta=0.05$, we would need $M=7.109\cdot 10^{8}$.
Taking $p=\frac{1}{10}$, but with complete information, we make a similar observation. For both algorithms, adding an inaccuracy factor $p=\frac{1}{10}$ decreases the performance with a factor $0.95-0.97$ on average. Taking into consideration the performance of the exact versions of the algorithms, we see, like before, a score confidently above the worst-case approximation ratio of $1-1/\operatorname{\text{e}}$, whereas our theoretical results required taking multiple samples for each query $(i,j)$ to compensate for the inaccuracy. Example 6.2 shows a computation of the theoretically required values.
Example 6.2.
Theorem 5.1 provided us with a constant factor to multiply the number queries with in order to guarantee a $(1-1/\operatorname{\text{e}})$-approximate solution with probability $\delta>0$. We compute $U=\lceil-2\frac{\log((nm)/\delta)}{\log(4p(1-p))}\rceil$, and taking again as example vtaiwan.uberx, with $m=197$ and $n=1921$. Then, for $\delta=0.05$, $U=32$. For our experiments, we used $U=1$.
Taking both $p=\frac{1}{10}$ and incomplete information, not meeting the theoretically required number repetitions for both scenarios, the performance takes a noticeable hit. These query algorithms perform at around $0.75-0.85$ of the exact versions of the algorithm, with a much bigger standard deviation. Remember, however, that we only queried a handful of statements per voter, and voters were possibly inaccurate, whereas the exact versions of the algorithms have complete access to all accurate information.
7 Conclusion
We studied approval-based committee elections with the goal of maximizing the Chamberlin-Courant score under incomplete or inaccurate information. The study of this type of multi-winner elections, and the focus on diversity, are motivated by applications in digital democracy for the support of online large-scale deliberation.
Analytically, the paper determines bounds on the query complexity for the above settings, in particular: for the incomplete information setting, we establish upper bounds for adaptations to our setting of known algorithms for the Max Cover problem (Theorems Theorem 4.10 and 4.4) and a lower bound (Theorem B.1) for any non-adaptive algorithm. For the inaccurate information setting, we establish a lower bound for optimal approximations with high probability (Theorem 5.2), as well as an upper bound using a greedy algorithm (Theorem 5.1). Empirically, we complement the worst-case results by testing the behavior of the algorithms on augmented real-world data from Polis. Even when voters respond, possibly inaccurately, to just a handful of statements, our algorithms perform well compared to the algorithms with complete and accurate access to each voters’ ballots.
A first direction for future work is to query voters adaptively, to reduce the amount of queries required. Second, we hope to combine both incomplete and inaccurate information theoretically. Third, we wish to develop a more comprehensive theory of diversity, beyond the reliance on the Chamberlin-Courant score. Fourth, we aim to strengthen our experimental section by including synthetic data. Fifth, we intend to explore the significance of the above results also to applications in active learning [31], with the aim of optimizing the querying of a group annotators in distributed data labeling tasks (see, e.g., [22]).
Acknowledgments.
@@
Appendix A Proofs of Section 4
Lemma 4.1.
Algorithm 1 is $(1-\frac{1}{\operatorname{\text{e}}})$-approximate for the maximum coverage problem, and runs in time $\mathcal{O}(kmn)$.
For the proof we use the following notation: for $i\leq k$, let $\operatorname{\textsc{gr}}_{i}$ denote the increase in Max Cover score the greedy algorithm attains in iteration $i$. Also for $i\leq k$, write $\operatorname{\textsc{gr}}(i)$ for the Max Cover score of the greedy solution after $i$ iterations, so $\operatorname{\textsc{gr}}(k)$ is the final score. We write $\operatorname{\textsc{opt}}(k)$ for the score an optimal solution, consisting of $k$ candidates, would have.
@@ Appendix A Proofs of Section 4
We start with the approximation ratio. Like [16, Lemma 1], we first prove that, for $\ell\leq k$, $\operatorname{\textsc{gr}}_{\ell}\geq(\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}(\ell-1))/k$. Fix an optimal solution. At the start of iteration $\ell$, the difference between the greedy score and the optimal score is $\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}(\ell-1)$. Thus, there are at least this many elements covered in the (fixed) optimal solution, that have not been covered yet by greedy up to that point (it could be more: the set selected by greedy at the start of iteration $\ell$, needn’t be a subset of the optimal solution). These $\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}(\ell-1)$ elements are covered by the $k$ sets in the optimal solution so, by the pigeon hole principle, there exists a set in the optimal solution, that is not currently selected by greedy, that will cover at least $(\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}(\ell-1))/k$ of these elements. Then, in iteration $\ell$, the greedy algorithm picks the set that covers the most additional elements, so these must be at least $(\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}(\ell-1))/k$ many. We use this relation to obtain an expression for $\operatorname{\textsc{gr}}(k)$ in terms of $\operatorname{\textsc{opt}}(k)$. For all $k$,
@@
Regarding the time complexity, the algorithm chooses $k$ candidates sequentially. For each of these $k$ steps, we have to go over all $m$ candidates to check which adds most to the current score. Per candidate, we have to check for each of the $n$ voters whether it is already represented (we can keep track of the represented voters in one set). This gives a time complexity in $\mathcal{O}(kmn)$. ∎
Remark A.1.
@@
We proceed by contraposition. Assume $\operatorname{\textsc{gr}}(k)<k$, which is equivalent to $\operatorname{\textsc{gr}}(k)\leq k-1$, so (by the pigeonhole principle) there exists $i\leq k$ such that $\operatorname{\textsc{gr}}_{i}=0$. Because $\operatorname{\textsc{gr}}_{i}\leq\operatorname{\textsc{gr}}_{i-1}$ for all $i\leq k$, we must have (at least) $\operatorname{\textsc{gr}}_{k}=0$; no candidate with a yet uncovered voter is left in iteration $k$, since otherwise the greedy algorithm would pick that over an empty candidate and then $\operatorname{\textsc{gr}}_{k}>0$. At this point, $\operatorname{\textsc{gr}}(k-1)\leq k-1$ and no uncovered voters are left, so $\sum_{i\in V}\mathds{1}(|A(i)|\geq 1)\leq k-1$. ∎
Lemma 4.3.
For any non-trivial instance, greedy-$\epsilon$ is $\left(\left(1-\frac{1}{\operatorname{\text{e}}}\right)\cdot\left(\frac{1}{1+\epsilon(1-\frac{1}{\operatorname{\text{e}}})}\right)\right)$-approximate.
For $i\leq k$, define $\operatorname{\textsc{gr}}\text{-}\epsilon_{i}$ and $\operatorname{\textsc{gr}}\text{-}\epsilon(i)$ similarly to before.

### v2

@@ Introduction
Let $|c_{j}|\geq 1\ \forall j$, $\delta>0$, $\gamma\in(0,1)$, $m\geq t>k\geq 3$. Fix $\beta=C_{2}\frac{1-\gamma}{\gamma k\log k}$ for some constant $C_{2}$. Then, w.p. at least $1-\delta$, Algorithm 4 is $(1-1/e)\gamma$-approximate for Max Cover with query complexity $O\left(\left(\frac{\gamma k\log k}{1-\gamma}\right)^{3}m\log\left(\frac{m\gamma k^{2}\log k}{\delta(1-\gamma)}\right)\right)\in O_{\delta,\gamma,k}(m\log m)$.
Like with greedy-incomplete, the query complexity does not depend on $n$, and is asymptotically much smaller than querying the entire ballot $nm$. Unlike before, there is sizeable constant overhead increasing with $k$ and $\gamma$, and decreasing with $\delta$. Especially the dependence in $k$ is much stronger than before. For the same example, $m=1000,\,\gamma=0.85,\,\delta=0.05,\,k=8$ and $t=20$, Theorem 6 requires $9.52180\cdot 10^{11}$ queries. Only for $n>9.52180\cdot 10^{8}$ this is smaller than $nm$.
Inaccurate information
In the inaccurate information setting, voters do hand in complete ballots, but the voters are inaccurate in reporting their true approvals: we assume each query $A(i,j)$ is incorrect with a probability $p$. We model this by defining, for $p\in(0,\frac{1}{2})$, a $p$-inaccurate query as $A_{p}(i,j):=A(i,j)\oplus X$, where $X\sim\text{Bernoulli}(p)$ [^4] This is essentially the classification noise model from PAC learning [17] applied to our setting.. Samples of $X$ are independent between voters, between candidates, and between consecutive draws of the same query. If $p=0$, we would have accurate information, and with $nm$ queries we would find a $(1-1/\operatorname{\text{e}})$-approximate solution (e.g. using the greedy algorithm). When answers may be corrupted, we can simply pose every question multiple times to compensate for the uncertainty. We do this to obtain an upper bound on the number of queries required to still acquire a $(1-1/\operatorname{\text{e}})$-approximate solution with high probability, adapting Algorithm 1, attaining Algorithm 6 (see Appendix D). Moreover, we provide a matching lower bound (up to log-factors). Both results hold for the problem of optimizing over a matroid constraint.
Theorem 7.
@@
Let $p\in(0,\frac{1}{2})$, $\delta>0$, $n,m\in\mathbb{N}$. Then there exists an algorithm that is (11/e)(1-1/e)-approximate for Max Cover in the $p$-inaccurate query model w.p. at least $1-\delta$ and with query complexity $O(nm\log(nm/\delta))$.
Theorem 8.
Let $p\in(0,\frac{1}{2})$, $\delta>0$, $n,m\in\mathbb{N}$. Then any algorithm that is (11/e)(1-1/e)-approximate for Max Cover in the $p$-inaccurate query model w.p. at least $1-\delta$ has expected query complexity $\Omega(nm\log(1/\delta))$.
The two bounds match (up to log-factors). Unlike before, we access the entire ballot, making the bound depend on $n$.
Experiments
Motivation
We run our algorithms on real-life data and synthetic data with realistic structure. [^5] All code and data is available at https://github.com/martijnmartijnmartijnmartijn/Diverse-Committees-with-Incomplete-or-Inaccurate-Approval-Ballots First, we establish empirically that the greedy and local search algorithms with complete information (Algorithm 1 and 3) consistently achieve a higher Max Cover score than the standard approval voting, electing the $k$ most approved candidates, justifying the study of our algorithms in the first place. Second, although we have proven that our querying algorithms can overcome inaccurate and incomplete voter responses with constant overhead, this constant overhead can still be very large. We show that in practice our algorithms beat the optimal approximation ratio, even when using much fewer queries than theoretically required, and even when we combine inaccurate and incomplete information (a setting that we did not study theoretically).
Data
We test our algorithms on 18 open-use datasets provided by Polis [^6] https://github.com/compdemocracy/openData. We did multiple pre-processing steps, most importantly completing the Polis data. As anticipated, participating voters typically vote on only a small portion of the statements, so we obtain a sparse comment-participant-matrix. Since these discussions have been completed already, we are not able to query voters anymore, so, for the sake of running our querying algorithms, we need to artificially complete the data. For details on this and other pre-processing steps we implemented, we refer to Appendix E.
To obtain more data beyond the 18 Polis datasets, we generate synthetic data that are structurally similar to the Polis data. Specifically, we employ the $(q,\phi)$-resampling model to sample approval elections, introduced by [33]. In this model, $q\in[0,1]$ represents the fraction of approvals and $\phi\in[0,1]$ represents the spread of approvals, so that $\phi=0$ means all voters are identical and approve the exact same $\lfloor q\cdot m\rfloor$ candidates, while $\phi=1$ means each candidate is approved with probability $q$ independently so that the spread of approvals is maximal. After pre-processing the Polis data, we find $q=0.0891$ and $\phi=0.693$. Appendix E contains an explanation of how to arrive at these values. We sampled 100 elections according to the $(0.0891,0.693)$-resampling model with $n=1000$ and $m=400$, as $m/n=0.4$ on average for the Polis datasets, and the average number of voters is roughly $1000$.
Figure 1: Max Cover score of greedy-incomplete (Algorithm 2) and ls-incomplete (Algorithm 4) on the 18 Polis and 100 synthetic data sets for $k=8$. The algorithms are configured so that each voter is expected to be queried only $M=1,2,3,4,5$ times. This was done for $p=0$ (accurate responses) and $p=0.1$ (inaccurate responses). On each of the 118 data sets, we ran 50 random trials of the algorithms, plotting the mean and standard deviation. All scores shown are relative to the complete and accurate information setting.
Results
approval local search greedy
Polis $0.77\pm 0.071$ $0.80\pm 0.074$ $0.82\pm 0.085$
Synthetic $8\pm 0.0040 0.99±0$$9\pm 0.0029$ $0.99\pm 0.0019$
@@ Results
Table 2: Mean and standard deviation of the Max Cover score of the algorithms greedy, local search (Algorithms 2 and 4) and approval voting with complete information on the 18 Polis datasets and 100 synthetic datasets. We do 20 random trials of local search.
To answer our first question, we ran greedy (Algorithm 1), local search-$\beta$ (Algorithm 3) [^7] We took $\beta$ as in Theorem 6 and write local search going forward and approval voting on the 118 data sets, including 20 random trials for each dataset for Algorithm 3, because of the random starting committee. Table 2 shows the mean Max Cover scores of the algorithms over all runs, separately for Polis data and synthetic data. On Polis data the mean score of all three algorithms is around 0.8, while on synthetic data the scores are nearly 1. greedy performs better than local search, which in turn performs better than approval voting. For Polis data, this difference in scores lies within the standard deviation. However, we stress that on no iteration of any Polis dataset, approval voting ‘beat’ the other two algorithms. The standard deviation thus reflects that different Polis datasets admit different Max Cover scores (across all three algorithms). On average over all Polis data, greedy and local search score higher than approval voting by a factor of 1.063 ($\sigma=0.038$) and by a factor of 1.045 ($\sigma=0.032$), respectively. On the synthetic data, greedy always produced a committee with a higher Max Cover score than approval voting, while local search did so on 95 out of 100 datasets (on 4 datasets it lost to approval voting on 1 out of 50 random trials; on 1 dataset it lost on 6 out of 50 trials), but the score was still within a factor of 0.996 of the score of approval voting. On average over all synthetic data, greedy and local search score higher than approval voting by a factor of 1.013 ($\sigma=0.004$) and by a factor of 1.008 ($\sigma=0.004$), respectively. We therefore conclude that our two algorithms, and especially greedy, consistently outperform approval voting by a decent factor.
@@
To answer our second question, we take the scores of greedy and local search from Table 2 as our baseline and inspect how close we can get to these scores when information is incomplete and/or inaccurate, using a realistic number of queries. For the incomplete information model, taking $t=20$, we write $M$ for the expected number of query sets of size $t$ that each voter is presented with. We then configure greedy-incomplete (Algorithm 2) and ls-incomplete (Algorithm 4) so that $M=1,2,3,4,5$, which is much lower than what would theoretically be required. We do 50 random trials of each algorithm (due to the random querying of voters). For the inaccurate information model, we ignore the theoretically required repetitions of the queries, and run the standard greedy and local search algorithms, but with $p=0.1$. Finally, we combine both of the above by running Algorithm 2 and Algorithm 4 with the above values of $M$ but for $p=0.1$. See figure 1 for the Max Cover scores obtained in these experiments.
@@
We see that both algorithms with accurate responses ($p=0$) obtain a score very close (0.85-0.95) to versions of the algorithms with complete information, even when $M=1$. With complete information, both algorithms obtained an average Max Cover score of around $0.8$. Multiplying by $0.9$ still yields a score above the worst-case approximation ratio of $1-1/\operatorname{\text{e}}\approx 0.63$, even when the optimal solution would have a Max Cover score of 1. Example 9 highlights the great performance witnessed in the experiments for $p=0$.
Example 9.
@@
Figure 1 shows that for $M=5$, the Max Cover score attained by ls-incomplete, on average across 18 Polis datasets with 50 runs per set, is approximately 0.95 times the score attained by local search. The dataset vtaiwan.uberx has $m=197$ and $n=1921$. For these values of $m$ and $n$, taking $\gamma=0.95$, $k=8$, $t=20$, and $\delta=0.05$, we would need $M=7.109\cdot 10^{8}$.
With $p=0.1$, but with complete information, the performance decreases by a factor of 0.95-0.97 on average, which yields a score confidently above the worst-case approximation ratio of $1-1/\operatorname{\text{e}}$. As a comparison: taking again dataset vtaiwan.uberx with $\delta=0.05$ would require repeating each query 32 times, but here we do not repeat queries at all.
Taking both $p=0.1$ and incomplete information, not meeting the theoretically required number of queries for either scenario, the performance takes a noticeable hit. The algorithms perform worse than the versions with complete information by a factor of 0.7-0.8 for $M=1$ increasing to 0.85-0.95 for $M=5$. This again implies that we are above the worst-case approximation ratio. Note that the standard deviation can become relatively large here, especially for the Polis data, which are more varied than the synthetic data.
Outlook
Our work is the first to address diversity in approval-based committee elections in the context of online civic participation platforms. Measuring diversity by the Chamberlin-Courant score, we prove diverse committees can be found by querying only a small fraction of the voters, even when responses may be inaccurate. This remains true in the presence of external diversity constraints on the committee, such as quota. Our algorithms match lower bounds on the query complexity (up to log-factors). We verify these theoretical results empirically on both real-life and synthetic data.
Our results open up several directions for future research. First, it would be interesting to combine the incomplete and inaccurate information models explicitly in theory. Second, it would be desirable to lift our results on inaccurate information to richer error models, e.g., where the error probability changes over time and/or between voters. Third, by viewing whom to query on what as an optimization problem—‘the comment routing problem’ by [29]—we hope to lower the query complexity further. Fourth, we intend to explore the significance of our results for applications in active learning [30], with the aim of optimizing the querying of a group of annotators in distributed data labelling tasks (see, e.g., [22]).
Acknowledgments
@@
Appendix A A   Proof of Theorem 2
We prove that greedy-incomplete (Algorithm 2) can get arbitrarily close to the optimal approximation ratio for Max Cover using $O(m\log m)$ queries to voters. This result was stated as Theorem 2. We start by giving a new proof that the regular greedy algorithm (Algorithm 1) achieves the optimal approximation ratio of $1-\frac{1}{\operatorname{\text{e}}}$, and we adapt this proof for Lemma 13, and consequently for the main result, Theorem 2.
Lemma 10.
Algorithm 1 is $(1-\frac{1}{\operatorname{\text{e}}})$-approximate for the maximum coverage problem, and runs in time $O(kmn)$.
For the proof we use the following notation: for $i\leq k$, let $\operatorname{\textsc{gr}}_{i}$ denote the increase in Max Cover score the greedy algorithm attains in iteration $i$, and $\operatorname{\textsc{gr}}(i)$ for the Max Cover score of the greedy solution after $i$ iterations. This means $\operatorname{\textsc{gr}}(k)$ is the score after termination. We write $\operatorname{\textsc{opt}}(k)$ for the score an optimal solution, consisting of $k$ candidates, would have.
@@ Appendix A A   Proof of Theorem 2
We start with the approximation ratio. Like [15, Lemma 1], we first prove that, for $\ell\leq k$, $\operatorname{\textsc{gr}}_{\ell}\geq(\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}(\ell-1))/k$. Fix an optimal solution. At the start of iteration $\ell$, the difference between the greedy score and the optimal score is $\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}(\ell-1)$. Thus, there are at least this many elements covered in the (fixed) optimal solution, that have not been covered yet by greedy up to that point (it could be more: the set selected by greedy at the start of iteration $\ell$, needn’t be a subset of the optimal solution). These $\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}(\ell-1)$ elements are covered by the $k$ sets in the optimal solution so, by the pigeon hole principle, there exists a set in the optimal solution, that is not currently selected by greedy, that will cover at least $(\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}(\ell-1))/k$ of these elements. Then, in iteration $\ell$, the greedy algorithm picks the set that covers the most additional elements, so these must be at least $(\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}(\ell-1))/k$ many. We use this relation to obtain an expression for $\operatorname{\textsc{gr}}(k)$ in terms of $\operatorname{\textsc{opt}}(k)$. For all $k$,
@@
Regarding the time complexity, the algorithm chooses $k$ candidates sequentially. For each of these $k$ steps, we have to go over all $m$ candidates to check which adds most to the current score. Per candidate, we have to check for each of the $n$ voters whether it is already represented (we can keep track of the represented voters in one set). This gives a time complexity in $O(kmn)$. ∎
Remark 11.
@@
We proceed by contraposition. Assume $\operatorname{\textsc{gr}}(k)<k$, which is equivalent to $\operatorname{\textsc{gr}}(k)\leq k-1$, so (by the pigeonhole principle) there exists $i\leq k$ such that $\operatorname{\textsc{gr}}_{i}=0$. Because $\operatorname{\textsc{gr}}_{i}\leq\operatorname{\textsc{gr}}_{i-1}$ for all $i\leq k$, we must have (at least) $\operatorname{\textsc{gr}}_{k}=0$; no candidate with a yet uncovered voter is left in iteration $k$, since otherwise the greedy algorithm would pick that over an empty candidate and then $\operatorname{\textsc{gr}}_{k}>0$. At this point, $\operatorname{\textsc{gr}}(k-1)\leq k-1$ and no uncovered voters are left, so $\sum_{v_{i}\in V}\mathds{1}_{\{|A(i)|\geq 1\}}(i)\leq k-1$, which means the instance is trivial. ∎
For the incomplete information setting we need to estimate approval information of the total population based on the information of only few voters. This can lead to errors. Write $\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon$ for any adaptation of Algorithm 1 that, in each iteration, chooses any candidate that yields the largest increase in score up to $\epsilon$; see Algorithm 5 for an example. We obtain the following approximation.
Lemma 13.
For any non-trivial instance, any greedy-$\epsilon$ algorithm is $\left(\left(1-\frac{1}{\operatorname{\text{e}}}\right)\cdot\left(\frac{1}{1+\epsilon(1-\frac{1}{\operatorname{\text{e}}})}\right)\right)$-approximate.
For $i\leq k$, define $\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon_{i}$ and $\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(i)$ similarly to before.